The supplement of an angle is 60 degrees less than twice the supplement of the complement of the angle. Find the measure of the complement.
Answer by ikleyn(52798)

.
The supplement of an angle is 60 degrees less than twice the supplement of the complement of the angle.
Find the measure of the complement.
~~~~~~~~~~~~~~~~~

Let x be  the angle  (which is called  "an angle"  in the problem's description).


Then the supplement is  (180-x) degrees,

and the complement is (90-x) degrees, while the supplement of the complement of the angle is  (180 - (90-x)) = 90+x degrees.


Next, write the equation as you read the text


    180-x = 2*(90+x) - 60.


Simplify and find x


    180-x = 180 + 2x - 60

       -x =       2x - 60

       60 =       2x + x

       60 =       3x

        x =       60/3 = 20.


Thus the angle is 20 degrees, and its complement (which is under the problem's question) is 90-20 = 70 degrees.    ANSWER

Solved and explained.

-------------

To see many other similar  (and different)  solved problems on supplementary and complementary angles,  look into the lessons
    - Angles basics
    - Vertical angles
    - HOW TO solve problems on supplementary, complementary or vertical angles - Examples
    - Solved problems on supplementary and complementary angles
in this site.

Consider these lessons as your handbook,  textbook,  guide,  tutorials and  (free of charge)  home teacher.